# Page curve entanglement dynamics in an analytically solvable model

### v2

@@ Abstract
The entanglement entropy of black holes is expected to follow the Page curve. After an initial linear increase with time the entanglement entropy should reach a maximum at the Page time and then decrease. This bending down of the Page curve and the apparent contradiction with Hawking’s semiclassical calculation from 1975 is at the center of the black hole information paradox. Motivated by this – from the point of view of non-equilibrium quantum many-body systems – unusual behavior of the entanglement entropy, this paper introduces an exactly solvable model of free fermions that explicitly shows such a Page curve: The entanglement entropy vanishes asymptotically for late times instead of saturating at a volume law. Physical observables like the particle current do not show any unusual behavior at the Page time and one can explicitly see how the semiclassical connection between particle current and entanglement generation breaks down.
@@
Motivated by this state of affairs this paper introduces a simple exactly solvable free fermion system plus environment model that shows Page curve like entanglement dynamics. This is interesting from the condensed matter point of view because – as explained above – such entanglement behavior is different from the generic picture Calabrese and Cardy (2005); Kim and Huse (2013); Ho and Abanin (2017). In fact, the dynamics will turn out to be quite intriguing after the equivalent of the Page time. In addition, the asymptotic quantum state in the environment has properties which cannot be generated by conventional quench protocols. With respect to black hole physics this paper makes no claim to contribute to the fundamental issue of the Hawking information paradox, although one can learn something about the difficulty in using non-rigorous arguments for a quantity as subtle as the entanglement entropy.
System plus bath quantum many-body models with random unitaries that generate a Page curve have previously been discussed in Refs. Liu and Vardhan (2021); Lau et al. (2022). A somewhat similar model to the one in this paper was used in Ref. Chen et al. (2020a), however, it yields discontinuous behavior at the Page time and not an actual bending down of the curve. Recent holographic calculations using the island formula have established that this framework desribes Page curve physics, though the entanglement entropy becomes constant instead of bending down at the Page time Chen et al. (2020b); Geng et al. (2021, 2022). The phenomenon of quantum distillation introduced in Ref. Heidrich-Meisner et al. (2009) leads to a decreasing entanglement entropy, but for different reasons than Page curve physics, which is the focus of this paper.
@@
Initially $m(0)=M$ and after a short transient one expects a constant current $\dot{m}(t)=I$, which can be thought of as the production of a particle-hole pairs at the boundary between system and environment: The holes travels into the system ${\cal S}$ and the particles into the environment ${\cal E}$. Semiclassically particle current and entanglement generation are proportional to one another Calabrese and Cardy (2005), so this leads to linear entanglement growth with a proportionality factor set by the particle current $I$. At late times the particles will be spread out approximately evenly throughout the whole system $\cal S$ and we therefore expect
$\lim_{t\rightarrow\infty}m(t)=\frac{M^{2}}{M+N}$ (8)
@@
$\Gamma_{k}(\epsilon)=\frac{2\pi\rho(\epsilon)\,g^{2}}{M+1}\,\sin^{2}\left(\frac{\pi\,k}{M+1}\right)\ .$ (22)
The reduced density operator of the system $\rho_{\rm sys}(t)$ is then the direct product of reduced density operators of the $M$ resonant level models
@@
with the dimensionless parameter $\tau=4\pi\rho g^{2}t/M$. Eqs. (2729) allow us to plot the entanglement entropy as a function of the fractional decay of the system, $S_{\rm ent}=S_{\rm ent}(m/M)$ such that the parameters $g$ and $\rho$ drop out. The resulting curves (shown in the next section) display the universal behavior in the limit $t_{\rm env}\gg t_{\rm sys}\gg g$, $N,M\rightarrow\infty$ while keeping $N\gg M^{2}$.
Entanglement entropy. Starting from the initial state (5) the system $\cal S$ starts to empty as can be seen in Fig. 3 obtained from the numerical solution of the Heisenberg equations of motion. Based on the discussion of (Page curve entanglement dynamics in an analytically solvable model) we therefore expect to see Page curve entanglement dynamics. Fig. 4 is obtained from the numerical solution following the Eisler-Peschel formalism described above and clearly exhibits this behavior: The entanglement entropy increases approximately linearly at early times, reaches its maximum at the Page time $t_{\rm P}\propto M$ and then decays again. The entanglement entropy at the Page time is proportional to the system size, $S_{\rm ent}(t_{\rm P})/M\approx 0.53\pm 0.02$. The proportionality factor is smaller than the Page value $\ln 2$ Page (1993), which is to be expected for a non-interacting system.
@@
Figure 4: Entanglement entropy as a function of time (top: linear scale, bottom: logarithmic scale). The Page curve behavior is clearly visible with the Page time $t_{\rm P}\propto M$ and $S_{\rm ent}(t_{\rm P})\propto M$. Parameters: $t_{\rm env}=4,t_{\rm sys}=1,g=0.5,N=10^{4}$.
Since the Page time depends on the coupling $g$ between system and environment, a universal way to depict the Page curve in our model is to plot the entanglement entropy as a function of the fraction of particles emitted into the environment, $1-m(t)/M$. The corresponding curves for different values of $g$ are shown in Fig. 5. Specifically, one can observe that the numerical results agree very well with the analytical universal curve from Eqs. (2729) (dashed line) for small values of $g$. The remaining deviations at late times are due to numerical limitations for the ratio $t_{\rm env}/t_{\rm sys}$ which one cannot choose much larger for $N=10^{4}$.
@@
This is depicted in Fig. 6 for the purity $S^{(2)}(t)$ and in Fig. 7 for the min-entropy $S^{(q=\infty)}(t)$. For larger values of $q$ the bending down at the Page time becomes more sudden with little indication of it at earlier times, see Fig. 7.
Figure 5: Entanglement entropy as a function of the fraction of particles that have been emitted into the environment for various system-environment couplings $g$ (3). The curves become universal in the weak-coupling limit $g\rightarrow 0$ and approach the analytical result from the disjoint resonant level model description (dashed line). Parameters: $t_{\rm env}=4,t_{\rm sys}=1,M=50,N=10^{4}$.
@@
Figure 7: Min-entropy $S^{({\rm min})}$ as a function of the fraction of particles that have been emitted into the environment for various system-environment couplings $g$. The dashed line is the analytical result from the disjoint resonant level model description. Parameters: $t_{\rm env}=4,t_{\rm sys}=1,M=50,N=10^{4}$.
@@
$\lim_{t\rightarrow\infty}\Delta H_{\rm env}^{2}(t)=\Delta H^{2}(t)=t_{\rm sys}^{2}\ .$ (35)
@@
Interpretation and generalizations. From the point of view of an observer in the system $\cal S$ the dynamical behavior of the entanglement or the energy variance is not surprising. Initially, system plus environment are in an unentangled product state (5). The process of emitting particles into the environment generates a complicated state in the system $\cal S$ that is entangled with the environment, but ultimately this process continues for so long that the system $\cal S$ is driven into a very small effective Hilbert space ${\cal H}^{\rm(eff)}_{\rm sys}(t\rightarrow\infty)$. Therefore entanglement and environment energy variance are forced to decrease again since the asymptotic small effective Hilbert space does not permit larger values. Fig. 9 gives a schematic picture of this process. Notice that in our model we have ${\rm dim}\,{\cal H}^{\rm(eff)}_{\rm sys}(t\rightarrow\infty)=1$ for $N\gg M^{2}$ and hence $S_{\rm ent}(t\rightarrow\infty)=0$. However, even for smaller ratios $N/M^{2}$ one can expect to see Page curve behavior in the sense that the entanglement entropy decreases after the Page time, albeit it will not decrease all the way to zero if ${\rm dim}\,{\cal H}^{\rm(eff)}_{\rm sys}(t\rightarrow\infty)>1$. Such finite size behavior is depicted in Fig. 10 and would be important for experimental realizations where one might not achieve $N/M^{2}\rightarrow 0$. Likewise the observations made here will carry over to interacting systems or higher dimensions since the mechanism in Fig. 9 is universal (however, then one is numerically limited to much smaller models).
@@
Figure 10: Entanglement entropy as a function of time for varying environment sizes $N$. The system size is $M=50$, other parameters $t_{\rm env}=t_{\rm sys}=1,g=0.65$. For this set of parameters the decrease of entropy after the Page time starts to set in for $N/M\approx 1.5$ and is clearly noticeable once $N/M\gtrsim 4$. The eventual increase of the entanglement entropy for even longer times (as the curves start to deviate from the $N=10^{4}$ curve) is due to the system not emptying out completely into the environment and particles being reflected at the right boundary of the environment.
@@
In contrast to such fine-tuned and highly sensitive scenarios with a decreasing entropy, models that follow Fig. 9 like the exactly solvable model analyzed in this paper show a robust decrease of the entanglement entropy after the Page time. Notice that at the Page time the entanglement entropy can be made arbitrarily large by increasing $M$. Even time-dependent perturbations to the system or the environment will not change the qualitative behavior of decreasing entanglement entropy after the Page time as long as the system $\cal S$ empties into the environment, i.e. as long as Fig. 9 remains applicable. An observer limited to the environment with no knowledge of the full initial state would conclude from such a robust decrease of the entanglement entropy that one is watching a movie running backwards. The dynamical buildup of long range entanglement between ${\cal H}_{\rm sys}$ and ${\cal H}_{\rm env}$ up to the Page time is ultimately responsible for the robust Page curve behavior: Particles emitted into the environment after the Page time carry entanglement across the boundary leading to the bending down of the Page curve. This behavior is robust as long as time evolution is unitary and one does not couple to additional external degrees of freedom.
Finally, it should be emphasized that one cannot make similarly general statements about a bending down of coarse-grained entropies based on Fig. 9 since this will depend on the specific coarse-graining procedure employed Almheiri et al. (2021).
@@
Summary. This paper introduced an exactly solvable model as an example for the general scenario depicted in Fig. 9 resulting in a Page curve. Unitary time evolution leads to an increase of the entanglement entropy until the Page time consistent with the semiclassical propotionality between particle current and entanglement generation (30). At the Page time the entanglement entropy can be arbitrarily large since it is proportional to the number of particles $M$ initially in the system ${\cal S}$. Beyond the Page time the entanglement entropy decreases again and this semiclassical picture breaks down. The scenario depicted in Fig. 9 generates entanglement dynamics different from the saturation at the volume law scenario commonly discussed in the literature without the need for any fine tuning. We have therefore constructed a solvable model that illustrates Polchinski’s burning piece of coal Polchinski (2016): The early photons are entangled with the remaining coal, but when the coal has burned completely the outgoing photons must be in a pure state (assuming the coal was initially in a pure state). While this is fundamentally different from a black hole due to the lack of an event horizon Polchinski (2016), one can make two observations in this exactly solvable model that have analogues in black hole physics: i) The semiclassical picture breaks down for a subtle quantity like the entanglement entropy once the system $\cal S$ in ${\cal H}_{\rm sys}$ has evolved to a sufficiently complex state at the Page time Akers et al. . ii) At early times there is little or no indication of the eventual bending down of the Page curve, see Figs. 4-7.

### v3

@@ Abstract
The entanglement entropy of black holes is expected to follow the Page curve. After an initial linear increase with time the entanglement entropy should reach a maximum at the Page time and then decrease. This paper introduces an exactly solvable model of free fermions that explicitly shows such a Page curve: The entanglement entropy vanishes asymptotically for late times instead of saturating at a volume law. The bending down of the Page curve is accompanied by a breakdown of the semiclassical connection between particle current and entanglement generation, a quantum phase transition in the entanglement Hamiltonian and non-analytic behavior of the $q\rightarrow\infty$ Renyi entropy. These observations are expected to hold for a larger class of systems beyond the exactly solvable model analyzed here.
@@
Motivated by this state of affairs this paper introduces a simple exactly solvable free fermion system plus environment model that shows Page curve like entanglement dynamics. This is interesting from the condensed matter point of view because – as explained above – such entanglement behavior is different from the generic picture Calabrese and Cardy (2005); Kim and Huse (2013); Ho and Abanin (2017). In fact, the dynamics turns out to be quite intriguing after the equivalent of the Page time defined as the maximum of the entanglement entropy curve. This exemplifies the difficulty of using non-rigorous arguments for a quantity as subtle as the entanglement entropy. Interestingly, the Page curve dynamics of our model is tied to a quantum phase transition of the entanglement Hamiltonian and non-analytic behavior of the $q\rightarrow\infty$ Renyi entropy at a critical time preceding the Page time. In addition, the asymptotic quantum state in the environment has properties which cannot be generated by conventional quench protocols. With respect to black hole physics this paper makes no claim to contribute to the fundamental issue of the Hawking information paradox, although there are some intriguing parallels as will be discussed below.
System plus bath quantum many-body models with random unitaries that generate a Page curve have previously been discussed in Refs. Liu and Vardhan (2021); Lau et al. (2022). A somewhat similar model to the one in this paper was used in Ref. Chen et al. (2020a), however, it yields discontinuous behavior at the Page time and not an actual bending down of the curve. Recent holographic calculations using the island formula have established that this framework describes Page curve physics, though the entanglement entropy becomes constant instead of bending down at the Page time Chen et al. (2020b); Geng et al. (2021, 2022). The phenomenon of quantum distillation introduced in Ref. Heidrich-Meisner et al. (2009) leads to a decreasing entanglement entropy, but for different reasons than Page curve physics, which is the focus of this paper.
@@
Initially $m(0)=M$ and after a short transient one expects a constant current $\dot{m}(t)=I$, which can be thought of as the production of a particle-hole pairs at the boundary between system and environment: The holes travels into the system ${\cal S}$ and the particles into the environment ${\cal E}$. Semiclassically particle current and entanglement generation are proportional to one another Calabrese and Cardy (2005), so this leads to linear entanglement growth with a proportionality factor set by the particle current $I$. At late times the particles will be spread out approximately evenly through system $\cal S$ and environment $\cal E$ yielding
$\lim_{t\rightarrow\infty}m(t)\propto\frac{M^{2}}{M+N}$ (8)
@@
$\Gamma_{k}(\epsilon)=\frac{2\pi\rho(\epsilon)\,g^{2}}{M+1}\,\sin^{2}\left(\frac{\pi\,k}{M+1}\right)$ (22)
where the operators $a^{\dagger}_{k,i},a_{k,i}$ are eigenoperators of the environment Hamiltonian (2). The reduced density operator of the system $\rho_{\rm sys}(t)$ is then the direct product of reduced density operators of the $M$ resonant level models
@@
with the dimensionless parameter
$\tau=4\pi\rho g^{2}t/M\ .$ (30)
Eqs. (2729) allow us to plot the entanglement entropy as a function of the fractional decay of the system, $S^{\rm(vN)}=S^{\rm(vN)}(m/M)$ such that the parameters $g$ and $\rho$ drop out. The resulting curves (shown in the next section) display the universal behavior in the limit $t_{\rm env}\gg t_{\rm sys}\gg g$, $N,M\rightarrow\infty$ while keeping $N\gg M^{2}$.
Entanglement entropy. Starting from the initial state (5) the system $\cal S$ starts to empty as can be seen in Fig. 3 obtained from the numerical solution of the Heisenberg equations of motion. Based on the discussion of (Page curve entanglement dynamics in an analytically solvable model) we therefore expect to see Page curve entanglement dynamics. Fig. 4 is obtained from the numerical solution following the Eisler-Peschel formalism described above and clearly exhibits this behavior: The entanglement entropy increases approximately linearly at early times, reaches its maximum at the Page time $t_{\rm P}\propto M$ and then decays again. The entanglement entropy at the Page time is proportional to the system size, $S^{\rm(vN)}(t_{\rm P})/M\approx 0.53\pm 0.02$. The proportionality factor is smaller than the Page value $\ln 2$ Page (1993), which is to be expected for a non-interacting system.
@@
Figure 4: Von Neumann entanglement entropy as a function of time (top: linear scale, bottom: logarithmic scale). The Page curve behavior is clearly visible with the Page time $t_{\rm P}\propto M$ and $S^{\rm(vN)}(t_{\rm P})\propto M$. Parameters: $t_{\rm env}=4,t_{\rm sys}=1,g=0.5,N=10^{4}$.
Since the Page time depends on the coupling $g$ between system and environment, a universal way to depict the Page curve in our model is to plot the entanglement entropy as a function of the fraction of particles emitted into the environment, $1-m(t)/M$. The corresponding curves for different values of $g$ are shown in Fig. 5. Specifically, one can observe that the numerical results agree very well with the analytical universal curve from Eqs. (2729) (dashed line) for small values of $g$.
The remaining deviations at late times are due to the finiteness of the environment in the numerical simulations: Each eigenmode of the system effectively couples to $N\,\Gamma_{k}/t_{\rm env}$ modes in the environment, leading to a proportionality factor $(t_{\rm env}/g)^{2}$ in (8). This implies that the exponential decay (26) only persists up to a finite residual occupation scaling like $(t_{\rm env}/g)^{2}\times M/N$, which in turn shifts the finite-$N$ entanglement curves above the RLM result. Going to even larger values of $N$ decreases these small deviations, but becomes numerically expensive and does not provide insights which seem to justify this effort.
@@
This is depicted in Fig. 6 for the purity $S^{(2)}(t)$ and in Fig. 7 for the min-entropy $S^{({\rm min})}(t)=\lim_{q\rightarrow\infty}S^{(q)}(t)$. For $q\rightarrow 1$ one obtains the usual von Neumann entanglement entropy $S^{\rm(vN)}(t)=\lim_{q\rightarrow 1}S^{(q)}(t)$. Notice that the min-entropy is just determined by the largest eigenvalue of the Schmidt decomposition
$S^{({\rm min})}(t)=-\ln\lambda_{1}(t)$ (33)
with
$|\Psi(t)\rangle=\sum_{i=1}^{M}\sqrt{\lambda_{i}}\,|\phi_{{\rm sys},i}(t)\rangle\otimes|\Phi_{{\rm env},i}(t)\rangle$ (34)
and $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{M}$. Interestingly, the min-entropy becomes non-analytic at the time indicated by the arrows in Fig. 7. This will be discussed in detail in the next section.
Figure 5: Von Neumann entanglement entropy as a function of the fraction of particles that have been emitted into the environment for various system-environment couplings $g$ (3). The curves become universal in the weak-coupling limit $g\rightarrow 0$ and approach the analytical result from the disjoint resonant level model description (dashed line). Parameters: $t_{\rm env}=4,t_{\rm sys}=1,M=50,N=10^{4}$.
@@
Figure 7: Min-entropy $S^{({\rm min})}$ as a function of the fraction of particles that have been emitted into the environment for various system-environment couplings $g$. The dashed line is the analytical result from the disjoint resonant level model description. The arrows mark the critical time with non-analytic behavior of the min-entropy in the resp. curve. Parameters: $t_{\rm env}=4,t_{\rm sys}=1,M=50,N=10^{4}$.
Non-analytic behavior of $S^{({\rm min})}$. The non-analytic behavior of $S^{({\rm min})}$ at a critical time (or equivalently: at a critical filling) is difficult to discern from Fig. 7. However, it can easily be deduced from analytical considerations. Let us first analyze the disjoint resonant level model description. Eqs. (24) and (33) imply for large $M$
[…]
This function is plotted in Fig. 8. For $\tau\leq\tau^{*}=\ln 2$ one has $n_{k}(\tau)\geq 1/2$ for all $k\in[0,\pi]$ and therefore
$\frac{S^{(\rm min)}(\tau)}{M}=\frac{\tau}{2}\ .$ (36)
For $\tau>\tau^{*}$ one has instead $21-n_{k}(\tau)\geq 1/$ 2 for the interval $I_{\rm empty}$ of $k$-values centered around $\pi/2$ (lower panel of Fig. 8) because the decay rates $\Gamma_{k}$ are largest in the middle of the band,
$I_{\rm empty}=\{k\in[0,\pi]\>|\>n_{k}(\tau)<1/2\}\ .$ (37)
For $\tau>\tau^{*}$ one can write
$\frac{S^{(\rm min)}(\tau)}{M}=\frac{\tau}{2}-k(\tau)$ (38)
with
[…]
An expansion around $\tau=\tau^{*}+\delta\tau=\ln 2+\delta\tau$, $\delta\tau\geq 0$ yields
$k(\tau^{*}+\delta\tau)=k_{0}\,\delta\tau^{3/2}+O(\delta\tau^{5/2})$ (41)
with $k_{0}=8/3\pi\sqrt{\ln 2}$.
This implies that $S^{(\rm min)}(\tau)$ is non-analytic at $\tau=\tau^{*}$ as clearly visible in Fig. 8 in the sense that the second time derivative becomes discontinuous,
$\lim_{\tau\uparrow\tau^{*}}\frac{d^{2}}{d\tau^{2}}\,\left(\frac{S^{(\rm min)}(\tau)}{M}\right)$ $=$ 00 0
[…]
Figure 8: Upper panel: The black line shows the min-entropy $S^{({\rm min})}$ of the disjoint resonant level model description as a function of dimensionless time $\tau$ from (30). The yellow line is the analytic continuation based on the early time behavior according to (36). Lower panel: The dark region depicts values of the wave vector $k$ with $n_{k}(\tau)>1/2$, the light region values of $k$ with $n_{k}(\tau)<1/2$.
While studies of the entanglement entropy often focus on the von Neumann entanglement entropy, the min-entropy also has an important quantum information theoretic interpretation: $S^{({\rm min})}$ determines the one-shot distillable entanglement Brandao and Datta (2011) and gives the distance to the closest separable state
$|D(t)\rangle=|\phi_{\rm sys}(t)\rangle\otimes|\Phi_{\rm env}(t)\rangle$ (43)
in the sense of the Fubini-Study metric Lockhart and Steiner (2002): If one chooses $|D(t)\rangle$ such that $|\langle D(t)|\Psi(t)\rangle|^{2}$ is maximized then
$-\ln\,|\langle D(t)|\Psi(t)\rangle|^{2}=S^{\rm(min)}(t)\ .$ (44)
In the Schmidt decomposition (34) this means
$|D(t)\rangle=|\phi_{{\rm sys},1}\rangle\otimes|\Phi_{{\rm env},1}\rangle\ .$ (45)
So an increasing min-entropy of $|\Psi(t)\rangle$ as a function of time means a larger distance to the closest unentangled state $|D(t)\rangle$. For $\tau<\tau^{*}$ the closest separable state is the one with all resonant level modes occupied
$|\phi_{{\rm sys},1}(\tau)\rangle=\prod_{k=0}^{\pi}c^{\dagger}_{k}\,|\Omega_{\rm sys}\rangle\ ,$ (46)
which is actually time independent for $\tau<\tau^{*}$. For $\tau>\tau^{*}$ the closest separable state has the system component
$|\phi_{{\rm sys},1}(\tau)\rangle=\prod_{k\in I_{\rm occ}(\tau)}c^{\dagger}_{k}\,|\Omega_{\rm sys}\rangle$ (47)
with the interval $I_{\rm occ}(\tau)=\{k\in[0,\pi]\>|\>n_{k}(\tau)\geq 1/2\}$. This means that some of the modes of the closest unentangled state are now empty as depicted in the lower panel of Fig. 8. The sudden appearance of a different closest separable state at $\tau=\tau^{*}$ is responsible for the dynamic transition and non-analytic behavior of $S^{\rm(min)}(\tau)$ at $\tau=\tau^{*}$.
Quantum phase transition of the entanglement Hamiltonian. Another point of view of the dynamical transition at $\tau^{*}$ can be obtained from the entanglement Hamiltonian (also called the modular Hamiltonian) defined via
$\rho_{\rm sys}(t)=\exp\left(-H_{\rm ent}(t)\right)\ .$ (48)
Defining the partition function
$Z_{\rm ent}(\beta)={\rm Tr}_{{\cal H}_{\rm sys}}\exp\left(-\beta\,H_{\rm ent}\right)$ (49)
and
$F_{\rm ent}(\beta)=-\beta^{-1}\,\ln\,Z_{\rm ent}(\beta)$ (50)
one can immediately derive from (32)
$S^{(q)}=\frac{q}{q-1}\,F_{\rm ent}(q)\ .$ (51)
So the Renyi entropies can be thought of as being proportional to the free energy of the entanglement Hamiltonian at inverse temperature $\beta=q$. Therefore the min-entropy with $q\rightarrow\infty$ corresponds to the zero temperature limit and is just the ground state energy of the entanglement Hamiltonian
$S^{\rm(min)}(t)=E^{\rm GS}_{\rm ent}(t)\ .$ (52)
Hence non-analytic behavior of the min-entropy is equivalent to non-analytic behavior of the ground state energy of the entanglement Hamiltonian, that is a quantum phase transition. Notice that because of (48) one has $Z(1)=1$, which implies $E^{\rm GS}_{\rm ent}(t)\geq 0$. Also one needs to be careful when taking the limit $q\rightarrow 1$ in (51) in order to obtain the von Neumann entanglement entropy.
For the disjoint resonant level model it is straightforward to derive the entanglement Hamiltonian from (24),
$H_{\rm ent}(\tau)=\sum_{k}\nu_{k}(\tau)\,c^{\dagger}_{k}c_{k}+a(\tau)\ .$ (53)
Here the single particle energies are given by
$\nu_{k}(\tau)=\ln\left(e^{\tau\sin^{2}k}-1\right)$ (54)
and the constant
$a(\tau)=-\sum_{k}\ln\left(1-e^{-\tau\sin^{2}k}\right)\ .$ (55)
For $\tau<\tau^{*}$ one has $\nu_{k}(\tau)<0$ for all $k$. Therefore all modes are occupied and the ground state energy is
$E^{\rm GS}_{\rm ent}(\tau)=a(\tau)+\sum_{k}\nu_{k}(\tau)\ .$ (56)
Once $\tau>\tau^{*}$ there are unoccupied modes with $\nu_{k}(\tau)>0$,
$E^{\rm GS}_{\rm ent}(\tau)=a(\tau)+\sum_{k:\nu_{k}(\tau)<0}\nu_{k}(\tau)\ .$ (57)
One can easily verify that (56) agrees with (36) and that (57) agrees with (38) as expected from (52), so one has a quantum phase transition of the entanglement Hamiltonian at $\tau=\tau^{*}$. This is accompanied by non-analytic behavior of the particle number in the ground state of the entanglement Hamiltonian,
$m^{\rm GS}_{\rm ent}\stackrel{{\scriptstyle\rm def}}{{=}}\langle{\rm GS}_{\rm ent}\,|\,\sum_{k}c^{\dagger}_{k}\,c_{k}\,|\,{\rm GS}_{\rm ent}\rangle\ .$ (58)
One has
$m^{\rm GS}_{\rm ent}(\tau)=\sum_{k\in I_{\rm occ}(\tau)}1$ (59)
yielding
$\frac{m^{\rm GS}_{\rm ent}(\tau)}{M}=\left\{\begin{array}[]{ll}1&\quad\tau\leq\tau^{*}\\ \frac{2}{\pi}\,\arcsin\sqrt{\frac{\tau^{*}}{\tau}}&\quad\tau>\tau^{*}\end{array}\right.$ (60)
This expression just corresponds to the occupied region in the lower panel of Fig. 8. At the transition one finds
$\frac{m^{\rm GS}_{\rm ent}(\tau^{*}+\delta\tau)}{M}=1-\frac{2}{\pi\sqrt{\ln 2}}\,\delta\tau^{1/2}+O(\delta\tau^{3/2})$ (61)
for $\delta\tau>0$.
The entanglement Hamiltonian also allows one to identify the critical times in the dynamics away from the small $g$ limit, see Fig. 7. Following Ref. Peschel and Eisler (2009) the single-particle energies $\nu_{i}$ of the entanglement Hamiltonian are the eigenvalues of $h$ from (13), so the quantum phase transition of $H_{\rm ent}(t)$ is determined by the first eigenvalue crossing zero since this implies that $m^{\rm GS}_{\rm ent}(\tau)$ drops from $M$ to $M-1$. This is depicted in Fig. 9 and this analysis located the critical times/fillings in Fig. 7.
Figure 9: Single particle energies $\nu_{i}$ of the entanglement Hamiltonian as a function of time. The overall trend is that the single particle energies become larger for later times since this corresponds to the emptying of the system into the environment. The upper panel shows a zoomed in plot where one can more clearly see when the first eigenvalue crosses zero (marked by red arrow), corresponding to the quantum phase transition of $H_{\rm ent}(t)$. Parameters: $t_{\rm env}=4,t_{\rm sys}=1,g=0.5,M=50,N=10^{4}$.
Looking at the Page curve from the point of view of the entanglement Hamiltonian also leads to a better understanding of the behavior of the finite-$q$ Renyi entropies $S^{(q)}$. According to (51) these can be identified with the free energy at inverse temperature $\beta=q$. Since the quantum phase transition of $H_{\rm ent}(t)$ is due to level crossing between sectors with different particle numbers, non-analytic behavior only occurs at zero temperature, that is in the min-entropy. All the finite-$q$ Renyi entropies remain analytic at all times, with the non-analytic behavior at the critical time emerging in the zero temperature ($q\rightarrow\infty$) limit. So while the critical behavior of $S^{\rm(min)}$ does not extend to finite $q$ Chandran et al. (2014), the bending down of the finite-$q$ entanglement entropies can continuously be traced back to the quantum phase transition at $q=\infty$ and the non-analytic behavior of the ground state particle number (60).
@@
$\lim_{t\rightarrow\infty}\Delta H_{\rm env}^{2}(t)=\Delta H^{2}(t)=g^{2}\ .$ (65)
@@
Interpretation and generalizations. From the point of view of an observer in the system $\cal S$ the dynamical behavior of the entanglement or the energy variance is not surprising. Initially, system plus environment are in an unentangled product state (5). The process of emitting particles into the environment generates a complicated state in the system $\cal S$ that is entangled with the environment, but ultimately this process continues for so long that the system $\cal S$ is driven into a very small effective Hilbert space ${\cal H}^{\rm(eff)}_{\rm sys}(t\rightarrow\infty)$. Therefore entanglement and environment energy variance are forced to decrease again since the asymptotic small effective Hilbert space does not permit larger values. Fig. 11 gives a schematic picture of this process. Notice that in our model we have ${\rm dim}\,{\cal H}^{\rm(eff)}_{\rm sys}(t\rightarrow\infty)=1$ for $N\gg M^{2}$ and hence $S^{\rm(vN)}(t\rightarrow\infty)=0$. However, even for smaller ratios $N/M^{2}$ one can expect to see Page curve behavior in the sense that the entanglement entropy decreases after the Page time, albeit it will not decrease all the way to zero if ${\rm dim}\,{\cal H}^{\rm(eff)}_{\rm sys}(t\rightarrow\infty)>1$. Such finite size behavior is depicted in Fig. 12 and would be important for experimental realizations where one might not achieve $N/M^{2}\rightarrow 0$. Likewise the observations made here will carry over to interacting systems or higher dimensions since the mechanism in Fig. 11 is universal (however, then one is numerically limited to much smaller systems).
@@
Figure 12: Von Neumann entanglement entropy as a function of time for varying environment sizes $N$. The system size is $M=50$, other parameters $t_{\rm env}=t_{\rm sys}=1,g=0.65$. For this set of parameters the decrease of entropy after the Page time starts to set in for $N/M\approx 1.5$ and is clearly noticeable once $N/M\gtrsim 4$. The eventual increase of the entanglement entropy for even longer times (as the curves start to deviate from the $N=10^{4}$ curve) is due to the system not emptying out completely into the environment and particles being reflected at the right boundary of the environment.
@@
In contrast to such fine-tuned and highly sensitive scenarios with a decreasing entropy, models that follow Fig. 11 like the exactly solvable model analyzed in this paper show a robust decrease of the entanglement entropy after the Page time. Notice that at the Page time the entanglement entropy can be made arbitrarily large by increasing $M$. Even time-dependent perturbations to the system or the environment will not change the qualitative behavior of decreasing entanglement entropy after the Page time as long as the system $\cal S$ empties into the environment, i.e. as long as Fig. 11 remains applicable. An observer limited to the environment with no knowledge of the full initial state would conclude from such a robust decrease of the entanglement entropy that one is watching a movie running backwards. The dynamical buildup of long range entanglement between ${\cal H}_{\rm sys}$ and ${\cal H}_{\rm env}$ up to the Page time is ultimately responsible for the robust Page curve behavior: Particles emitted into the environment after the Page time carry entanglement across the boundary leading to the bending down of the Page curve. This behavior is robust as long as time evolution is unitary and one does not couple to additional external degrees of freedom. It should be emphasized that one cannot make similarly general statements about a bending down of coarse-grained entropies based on Fig. 11 since this will depend on the specific coarse-graining procedure employed Almheiri et al. (2021).
Interestingly, in our model the bending down of the von Neumann and all other Renyi entanglement entropies can be related to a quantum phase transition of the entanglement Hamiltonian accompanied by non-analytic behavior of the min-entropy, see (42) and Fig. 8. Therefore the increasing and decreasing part of the Page curve are separated by a critical point of $S^{\rm(min)}$ (strictly speaking, the critical point slightly precedes the Page time, so this is a qualitative statement about the overall behavior). Currently it is unclear whether this observation of a dynamical transition can be generalized to the other forcing mechanisms from Fig. 11, which would be very intriguing. At least for models with a conserved charge this seems very likely since the quantum phase transition of the entanglement Hamiltonian is driven by a level crossing between different sectors of the conserved charge, which is unaffected by interactions. More work will be required to explore these questions.
@@
Summary. This paper introduced an exactly solvable model as an example for the general scenario depicted in Fig. 11 resulting in a Page curve. Unitary time evolution leads to an increase of the entanglement entropy until the Page time consistent with the semiclassical propotionality between particle current and entanglement generation (31). At the Page time the entanglement entropy is proportional to the number of particles $M$ initially in the system ${\cal S}$. Beyond the Page time the entanglement entropy decreases again and the semiclassical picture breaks down. The scenario depicted in Fig. 11 generates entanglement dynamics different from the saturation at the volume law scenario commonly discussed in the literature without the need for any fine tuning.
We have therefore constructed a solvable model that illustrates Polchinski’s burning piece of coal Polchinski (2016): The early photons are entangled with the remaining coal, but when the coal has burned completely the outgoing photons must be in a pure state (assuming the coal was initially in a pure state). While this is fundamentally different from a black hole due to the lack of an event horizon Polchinski (2016), one can make two observations in this exactly solvable model that have analogues in black hole physics: i) The semiclassical picture (31) breaks down for a subtle quantity like the entanglement entropy once the system $\cal S$ in ${\cal H}_{\rm sys}$ has evolved to a sufficiently complex state at the Page time Akers et al. . ii) The $q\rightarrow\infty$ Renyi entropy $S^{\rm(min)}$ exhibits non-analytic behavior at a critical time which slightly precedes the Page time (Fig. 8). The sudden onset of a non-analytic contribution according to (38) shows a certain resemblance to the island formula and the resulting non-analytic behavior in holographic models Almheiri et al. (2019); Penington et al. (2022). In our model the non-analytic contribution can be traced back to the sudden emergence of a new closest (in the sense of the Fubini-Study metric) separable state (47) to the time evolved initial state. Whether there is an underlying connection between these observations in holography and our model is a question for future study. In any case the quantum phase transition of the entanglement Hamiltonian discussed above is driven by a new mechanism and deserves further study from the quantum many-body point of view, especially regarding generalizations to interacting models with or without conserved charges and the implications for the finite-$q$ Renyi entropies.